# Repetition Improves Language Model Embeddings

### v1

@@ Appendix D Additional Finetuning Results
Validation Dataset Classification Pair Classification Clustering Retrieval STS Reranking Average
Classical
Classification 59.20 73.80 24.16 20.57 58.59 54.54 46.79
Pair Classification 58.73 71.40 24.32 20.39 59.00 54.42 46.64
Clustering 58.23 72.62 23.90 18.64 56.68 54.82 45.37
Retrieval 58.21 73.87 23.85 20.35 56.97 54.44 45.88
STS 58.31 44.03 13.07 2.63 38.95 46.77 34.63
Reranking 58.13 71.77 24.20 20.23 58.59 54.89 46.43
Echo
Classification 64.50 74.65 25.93 22.52 73.81 59.41 55.57
Pair Classification 64.15 75.93 22.25 18.35 72.75 58.47 54.15
Clustering 61.54 71.04 26.32 15.88 68.18 60.27 51.81
Retrieval 64.06 75.26 27.02 23.61 72.40 60.00 55.07
STS 64.50 74.65 25.93 22.52 73.81 59.41 55.57
Reranking 64.15 75.93 22.25 18.35 72.75 58.47 54.15
Summarization
Classification 66.62 78.95 21.79 14.68 72.13 64.24 55.22
Pair Classification 66.62 78.95 21.79 14.68 72.13 64.24 55.22
Clustering 66.66 79.59 28.08 11.88 67.30 65.19 53.43
Retrieval 66.01 81.82 26.48 19.13 70.13 66.24 54.96
STS 66.01 81.82 26.48 19.13 70.13 66.24 54.96
Reranking 63.19 75.22 26.09 20.52 65.98 59.05 51.55
Table 8: Scores for additional zero-shot validation datasets on Mistral-7B.
Validation Dataset Classification Pair Classification Clustering Retrieval STS Reranking Average
Classical
Classification 57.59 68.65 23.72 18.06 57.19 54.59 45.14
Pair Classification 57.56 70.18 23.51 18.54 58.24 54.40 45.79
Clustering 57.14 69.91 23.35 16.98 57.66 55.38 45.25
Retrieval 56.61 68.46 23.22 18.63 56.49 53.26 44.65
STS 57.56 70.18 23.51 18.54 58.24 54.40 45.79
Reranking 56.65 66.54 22.46 10.48 55.97 54.44 42.98
Echo
Classification 62.24 67.96 23.60 14.33 65.79 55.44 49.85
Pair Classification 63.42 72.52 21.11 17.35 68.16 54.98 51.47
Clustering 60.12 66.74 23.45 11.60 64.45 56.31 48.75
Retrieval 61.64 66.29 25.11 16.12 66.18 56.35 50.26
STS 63.15 68.74 23.65 16.38 69.37 57.75 51.96
Reranking 62.30 74.23 24.69 18.17 65.07 56.76 50.51
Summarization
Classification 63.96 77.93 21.89 15.93 67.07 63.39 52.34
Pair Classification 63.96 77.93 21.89 15.93 67.07 63.39 52.34
Clustering 61.60 69.47 24.44 5.28 57.53 57.62 45.85
Retrieval 64.90 78.74 26.63 15.59 70.15 65.43 54.02
STS 64.90 78.74 26.63 15.59 70.15 65.43 54.02
Reranking 60.54 69.73 26.40 15.82 61.60 58.80 47.83
Table 9: Scores for additional zero-shot validation datasets on LLaMa-2-7B.
Validation Dataset Classification Pair Classification Clustering Retrieval STS Reranking Average
Classical
Classification 58.24 71.65 23.91 21.79 58.74 56.37 46.66
Pair Classification 58.10 73.30 23.01 16.97 57.83 56.17 45.52
Clustering 58.61 67.47 23.30 15.51 57.93 56.86 45.05
Retrieval 58.50 65.06 24.22 18.92 57.47 56.38 45.15
STS 58.24 71.65 23.91 21.79 58.74 56.37 46.66
Reranking 58.61 67.47 23.30 15.51 57.93 56.86 45.05
Echo
Classification 64.15 74.22 25.02 27.58 70.81 61.43 55.02
Pair Classification 64.57 77.63 22.56 24.08 69.85 59.89 53.55
Clustering 63.26 73.50 25.10 27.48 69.04 61.81 54.32
Retrieval 64.65 74.57 25.72 26.58 72.20 62.68 55.60
STS 63.16 75.98 24.08 27.56 71.00 61.84 54.85
Reranking 62.90 70.58 25.53 22.11 68.82 62.38 53.02
Summarization
Classification 66.02 79.06 26.47 22.20 67.91 64.90 54.52
Pair Classification 66.02 79.06 26.47 22.20 67.91 64.90 54.52
Clustering 63.84 71.98 21.99 7.48 56.96 59.41 46.50
Retrieval 66.02 79.06 26.47 22.20 67.91 64.90 54.52
STS 66.02 79.06 26.47 22.20 67.91 64.90 54.52
Reranking 61.19 69.63 26.38 19.62 60.76 62.79 48.36
Table 10: Scores for additional zero-shot validation datasets on LLaMa-2-13B.
Dataset Classical Echo Summarization
FiQA2018 (retrieval) 7.89 12.74 12.43
SCIDOCS (retrieval) 3.60 4.88 9.97
SciFact (retrieval) 45.39 49.36 29.90
NFCorpus (retrieval) 12.07 16.57 17.51
TwitterSemEval20. (pair_classification) 47.81 62.49 59.79
TwitterURLCorpus (pair_classification) 73.87 75.26 81.82
ImdbClassificati. (classification) 72.50 72.02 82.78
AmazonReviewsCla. (classification) 37.09 40.72 45.58
TweetSentimentEx. (classification) 53.70 58.76 61.74
MTOPDomainClassi. (classification) 83.85 92.71 90.72
TwentyNewsgroups. (clustering) 20.84 29.48 30.11
BiorxivClusterin. (clustering) 23.47 27.61 27.21
MedrxivClusterin. (clustering) 24.23 26.42 25.75
StackOverflowDup. (reranking) 35.85 42.71 40.32
AskUbuntuDupQues. (reranking) 49.49 54.09 57.17
SciDocsRR (reranking) 59.38 65.91 75.30
BIOSSES (sts) 59.05 78.19 66.06
STS12 (sts) 42.01 58.43 64.62
STS13 (sts) 59.66 78.53 78.45
STS14 (sts) 50.69 68.42 71.00
STS15 (sts) 61.81 78.82 78.29
STS16 (sts) 57.03 77.52 77.40
STS17 (sts) 68.08 82.14 78.80
STS22 (sts) 61.23 57.60 47.07
STSBenchmark (sts) 47.55 73.85 77.39
SICK-R (sts) 53.19 71.95 69.48
Average 45.88 55.07 54.96
Table 11: Evaluation of all MTEB datasets for zero-shot for Mistral-7B.
Dataset Classical Echo Summarization
FiQA2018 (retrieval) 6.48 12.38 9.00
SCIDOCS (retrieval) 3.72 4.38 8.33
SciFact (retrieval) 42.18 30.61 23.01
NFCorpus (retrieval) 10.01 13.38 15.43
TwitterSemEval20. (pair_classification) 44.11 54.66 54.27
TwitterURLCorpus (pair_classification) 68.46 66.29 78.74
ImdbClassificati. (classification) 71.65 73.11 85.83
AmazonReviewsCla. (classification) 36.16 40.68 44.77
TweetSentimentEx. (classification) 52.04 54.85 59.96
MTOPDomainClassi. (classification) 81.63 89.38 89.97
TwentyNewsgroups. (clustering) 15.88 23.42 32.28
BiorxivClusterin. (clustering) 23.13 25.92 27.79
MedrxivClusterin. (clustering) 23.31 24.30 25.48
StackOverflowDup. (reranking) 35.57 40.82 35.63
AskUbuntuDupQues. (reranking) 48.51 51.42 56.09
SciDocsRR (reranking) 58.01 61.29 74.76
BIOSSES (sts) 65.31 71.96 68.04
STS12 (sts) 41.84 52.40 60.20
STS13 (sts) 58.43 72.40 76.31
STS14 (sts) 49.21 61.24 68.73
STS15 (sts) 60.03 72.67 75.59
STS16 (sts) 56.40 73.51 76.71
STS17 (sts) 62.31 71.87 79.38
STS22 (sts) 59.48 55.21 55.69
STSBenchmark (sts) 49.45 65.73 76.42
SICK-R (sts) 55.35 64.39 70.69
Average 44.65 50.26 54.02
Table 12: Evaluation of all MTEB datasets for zero-shot for LLaMa-2-7B.
Dataset Classical Echo Summarization
FiQA2018 (retrieval) 8.31 18.07 9.43
SCIDOCS (retrieval) 4.87 7.56 10.38
SciFact (retrieval) 41.64 50.55 40.19
NFCorpus (retrieval) 10.26 21.63 16.02
TwitterSemEval20. (pair_classification) 42.43 62.85 59.55
TwitterURLCorpus (pair_classification) 65.06 74.57 79.06
ImdbClassificati. (classification) 71.82 75.44 91.86
AmazonReviewsCla. (classification) 37.88 43.25 50.60
TweetSentimentEx. (classification) 52.95 58.18 59.93
MTOPDomainClassi. (classification) 84.67 92.52 87.51
TwentyNewsgroups. (clustering) 17.21 25.98 32.08
BiorxivClusterin. (clustering) 24.95 26.75 28.30
MedrxivClusterin. (clustering) 23.49 24.70 24.64
StackOverflowDup. (reranking) 37.24 44.86 38.44
AskUbuntuDupQues. (reranking) 50.74 55.21 54.15
SciDocsRR (reranking) 62.03 70.15 75.65
BIOSSES (sts) 63.26 77.60 69.33
STS12 (sts) 51.80 59.36 51.17
STS13 (sts) 61.59 79.01 76.08
STS14 (sts) 49.69 69.75 66.62
STS15 (sts) 58.48 79.86 73.75
STS16 (sts) 53.18 76.75 77.40
STS17 (sts) 65.10 80.41 75.88
STS22 (sts) 59.00 56.84 49.23
STSBenchmark (sts) 44.80 71.31 75.17
SICK-R (sts) 55.13 70.27 71.70
Average 45.15 55.60 54.52
Table 13: Evaluation of all MTEB datasets for zero-shot for LLaMa-2-13B.
@@
Model Average Clas. Clus. Pair Clas. Rera. Retr. STS Summ.
Classical (w/ instruct., mean) 62.96 76.26 42.68 86.31 57.58 53.75 81.53 30.19
Classical (w/ instruct., last) 63.98 76.57 45.78 86.37 56.71 54.87 82.03 31.02
Echo (w/ instruct., mean) 64.22 77.00 44.94 87.73 58.30 55.11 82.52 29.46
Echo (w/ instruct., last) 64.68 77.43 46.32 87.34 58.14 55.52 82.56 30.73
Classical (w/out instruct., mean) 62.19 75.23 41.79 85.24 56.31 53.24 80.97 30.64
Classical (w/out instruct., last) 62.37 75.01 42.70 85.69 56.64 53.29 80.92 30.91
Echo (w/out instruct., mean) 63.28 75.26 42.93 86.95 57.05 55.65 81.40 30.62
Echo (w/out instruct., last) 62.80 75.30 42.94 86.31 57.31 54.18 80.92 31.00
Classical (w/ instruct., mean, step 280) 63.19 76.18 42.99 85.44 57.63 53.96 82.53 29.94
Classical (w/ instruct., last, step 280) 63.87 76.54 46.22 86.70 57.79 53.73 82.22 30.13
Echo (w/ instruct., mean, step 280) 64.04 76.84 45.76 87.72 59.33 53.55 82.64 30.33
Echo (w/ instruct., last, step 280) 64.50 76.41 46.70 87.17 59.10 54.84 82.98 31.09
Table 15: Additional ablations for finetuning.
Dataset Repetition (last) Repetition (mean) Classical (last) Classical (mean) Bidirectional (last)

### v2

@@ Appendix C Echo Embeddings: Additional Information
Appendix D Additional Zero-shot Results
Recent literature suggests that the performance of language models on zero-shot tasks can be highly variable depending on the exact wording and template of the prompts (Sclar et al., 2023). Thus, in order to test the degree to which the exact prompt matters for each of the embedding strategies we consider, we perform prompt randomization where we sample prompts by randomizing the exact wording, punctuation, and capitalization of the prompt.
Prompt sampling procedure.
  1. 1.
    Choose an instruction. For classical embeddings, we choose from {Write, Say, Complete, Explain}. For echo embeddings, we choose from {Repeat, Rewrite, Rephrase, Fill in the blank}. For PromptEOL, we choose from {Summarize, Categorize, Understand, Analyze}.
  2. 2.
    Choose a wording for the instruction. For example, if we chose “Say” as the instruction, then we would choose from {Say a sentence, Say a paragraph, Say something, Say a response, Say a query, Say a prompt}. For PromptEOL, we also choose a second part of the wording, as the PromptEOL strategy requires that the summary be in one word: {in one word, with a single word, succinctly with one word, in a unique one-word way, in a single word, in a word}.
  3. 3.
    Choose a separator, which include colons, commas, newlines.
  4. 4.
  5. 5.
    Classical prompts have the form: “{instruction} {separator} {prefix} SS”.
  6. 6.
[…]
  7. 7.
    PromptEOL prompts have the form: “{instruction0} {separator} {prefix} $S$ {instruction1} {separator}”.
Classical embeddings.
For classical, we choose the prompts:
Write a sentence I] $S$
Write a prompt!
(I) $S$
Write some text
PROMPT-$S$
SAY A PARAGRAPH — SENTENCE 0] $S$
Say a query QUERY: $S$
Say a sentence!
[A] $S$
COMPLETE THE PROMPT Text (1) $S$
COMPLETE THE QUERY text 0) $S$
Complete the query SENTENCE 0) $S$
Complete the sentence:-$S$
Explain a query text 0 $S$
Explain a prompt — Sentence 1¿ $S$
EXPLAIN A SENTENCE Prompt (1) $S$
Echo embeddings.
For echo, we choose the prompts:
Repeat The Paragraph.
query 1) $S$.
query 2) $S^{\prime}$
Repeat the response.
1) $S$.
AGAIN 2) $S^{\prime}$
REPEAT THE SENTENCE :: PROMPT
$S$ :: RESPONSE
S’
Rewrite the query — QUERY (A) $S$ — (B) $S^{\prime}$
Rewrite the text. SENTENCE A) $S$. B) $S^{\prime}$
Rewrite the response — query A] $S$ — query B] $S^{\prime}$
Rephrase the sentence:@$S$:Again@S’
Rephrase The Sentence! Text ¡¿ SS! Answer ¡¿ $S^{\prime}$
REPHRASE THE QUERY Sentence a) $S$ Answer b) $S^{\prime}$
Fill in the blank in the prompt: Query a) $S$: Query b) $S^{\prime}$
FILL IN THE BLANK IN THE RESPONSE — Sentence A) $S$ — Sentence B) $S^{\prime}$
Fill in the blank in the paragraph.
Text — $S$.
Response — $S^{\prime}$
PromptEOL embeddings.
For PromptEOL, we use the prompts:
SUMMARIZE THE QUERY.
Prompt: $S$’IN A WORD.
Summarize the sentence!
PROMPT ¡1¿ $S$’Succinctly With One Word!
SUMMARIZE THE PARAGRAPH. PROMPT (0) $S$’IN A WORD.
CATEGORIZE THE PROMPT query
$S$ With a single word
Categorize the query — prompt [1] $S$ in a word —
CATEGORIZE THE SENTENCE.
Prompt ¡1¿ $S$ IN A WORD.
Understand the sentence
@$S$ In a single word
Understand The Prompt:QUERY [0] $S$ in a single word:
UNDERSTAND THE PARAGRAPH:Text I] $S$ Succinctly with one word:
Analyze the sentence.
Sentence $S$ In A Unique One-word Way.
Analyze the response! query a¿ $S$ IN A UNIQUE ONE-WORD WAY!
Analyze The Prompt
Sentence a¿ $S$ In a unique one-word way
Subset of MTEB for zero-shot sensitivity evaluation.
Given that we evaluate on hundreds of prompts, we choose a subset of MTEB to reflect each of the MTEB categories but with a substantially reduced computational requirement for evaluation. We evalaute on the following subset of MTEB: FiQA2018, SCIDOCS, SciFact, NFCorpus, TwitterSemEval2015, TwitterURLCorpus, ImdbClassification, AmazonReviewsClassification, TweetSentimentExtractionClassification, MTOPDomainClassification, TwentyNewsgroupsClustering, BiorxivClusteringS2S, MedrxivClusteringS2S, StackOverflowDupQuestions, AskUbuntuDupQuestions, SciDocsRR, BIOSSES, STS12, STS13, STS14, STS15, STS16, STS17, STS22, STSBenchmark, and SICK-R. Note that this excludes the largest retrieval datasets in MTEB, on which we find PromptEOL is most likely to fail.
We plot the sensitivity of classical, repetition, and PromptEOL to different choices of prompts for different models in Figures 78, and 9. We also extend to plotting on all tested datasets individually in Figures 1011, and 12. We observe that PromptEOL is highly sensitive to the exact prompt used. However, neither classical nor echo were particularly sensitive. Consistently, mean token pooling outperformed last token pooling by a large factor.
Appendix E Toy Verification
In Section 3, we demonstrate that classical embeddings overestimate the similarity between examples which are superficially similar based on tokens that appear early in the sequence. In this section, we investigate whether this failure mode generalizes to real data. We provide qualitative and quantitative evidence that this failure mode does, in fact, generalize to real data.
E.1 Qualitative comparison of classical and echo embeddings.
To build intuition that this applies to realistic data, we present the sentence pair from STSBenchmark, a sentence similarity task from MTEB, in which echo embeddings reduce error the most:
[…]
For these examples, the sentence pair $(x_{1},x_{2})$ on which echo embeddings improve error the most has exactly the property we identify as a failure mode for classical embeddings: the sentence is superficially similar for the first few tokens. On the other hand $(y_{1},y_{2})$ does not have this property.
E.2 Quantitative evaluation of the failure mode.
In order to test this hypothesis quantitatively, we design an experiment to identify real examples that are superficially similar in their early tokens, and examples that do not have this property. We compare the behavior of classical embeddings on these datapoints.
To conduct this experiment, we extract a set of examples from the STS datasets included in the MTEB benchmark in which the first half of the sentence is similar, and measure the degree to which the similarity is overestimated. We also select points which are similar in the second half of the sentence, and measure the degree to which similarity is overestimated. By comparing the degree to which sentences which are similar in the first half are overestimated in similarity, and the degree to which sentences which are similar in the second half are overestimated, then we can identify if classical embeddings overestimate similarity in specifically sentences which are similar in the first half. Thus, under our hypothesis, we expect that, for classical embeddings, sentences which are similar in the first half are overestimated in similarity more than sentences that are similar in their second half. On the other hand, we expect that, for echo embeddings, the degree to which similarity is over- or underestimated is independent of whether the sentences are similar in the first or second half of the sentence.
Measuring sentence similarity estimation error.
Results.
Qualitative examples.
E.3 Full zero-shot
Here we present omitted results from the main paper. First, the role of scale and base model is shown in Table 12
Table 12: Role of scale and base model in the zero-shot setting: We study the role of scale and base model using a different 7B parameter model: LLaMA-7B and a 1.3B parameter model: S-LLaMA-1.3B.
Categories $\longrightarrow$ Clas. Clus. P. Cls. Rera. Retr. STS Su. Average
# of datasets 12 11 3 4 15 10 1 56
Echo 1.3B 64.85 30.62 47.30 42.35 13.23 61.68 23.80 40.45
Classical 1.3B 59.41 27.50 41.01 38.09 8.53 47.91 23.36 34.31
PromptEOL 1.3B 68.86 24.52 36.52 43.81 8.34 56.64 27.36 37.50
Echo 7B 70.75 30.68 65.15 45.72 18.03 70.19 29.81 45.84
Classical 7B 63.54 31.99 53.78 42.82 15.07 55.69 26.21 40.29
PromptEOL 7B 74.02 26.93 58.04 47.48 10.11 66.94 29.71 42.84
@@ E.3 Full zero-shot
Appendix F Additional Finetuning Results
Table 13: MTEB Leaderboard: We compare to other recent models on MTEB. For a more complete list, please visit: https://huggingface.co/spaces/mteb/leaderboard.
Categories $\longrightarrow$ Clas. Clus. P. Cls. Rera. Retr. STS Su. Average
# of datasets 12 11 3 4 15 10 1 56
Main fine-tuning results trained with publicly available data:
Echo (ours) 77.43 46.32 87.34 58.14 55.52 82.56 30.73 64.68
Echo + compute matched 77.39 46.27 87.49 58.08 55.07 83.18 30.73 64.66
Classical 76.57 45.78 86.37 56.71 54.87 82.03 31.02 63.98
Classical + bidirectional attn. 76.70 45.94 88.15 57.23 54.96 82.42 29.32 64.23
UAE-Large-V1 (MLM) 75.58 46.73 87.25 59.88 54.66 84.54 32.03 64.64
multilingual-e5-large (MLM) 77.56 47.10 86.19 58.58 52.47 84.78 30.39 64.41
bge-large-en-v1.5 (MLM) 75.97 46.08 87.12 60.03 54.29 83.11 31.61 64.23
udever-bloom-7b (autoregr.) 72.13 40.81 85.40 55.91 49.34 83.01 30.97 60.63
sgpt-5.8b (autoregr.) 68.13 40.34 82.00 56.56 50.25 78.10 31.46 58.93
Prior work trained with closed-source data (see Section 5.2):
e5-mistral-7b (autoregr.) 78.47 50.26 88.34 60.21 56.89 84.63 31.40 66.63
GritLM-7B (autoregr.) 79.46 50.61 87.16 60.49 57.41 83.35 30.37 66.76
Training Datasets.
GPUs.
Training a model takes approximately two days on four A100 GPUs. Evaluating a model on the MTEB benchmark can be completed in parallel in approximately two days with eight A100s.
Instructions for finetuning datasets.
Models on MTEB leaderboard.
We compare our implementation of classical and echo embeddings to state-of-the-art approaches on MTEB. Namely, we display results for UAE-Large-V1 (Li & Li, 2023a), multilingual-e5-large (Wang et al., 2024), bge-large-en-v1.5 (Xiao et al., 2023b), udever-bloom-7b (Zhang et al., 2023a), sgpt-5.8b (Muennighoff, 2022), e5-mistral-7b (concurrent work) (Wang et al., 2023).
Pooling.
We investigate the role of pooling for fine-tuning (Table 14). We find that last-token, with a trainable end-of-sentence token embedding, is optimal.
Table 14: Role of pooling in the fine-tuning setting: Average MTEB Score (56 datasets) for different pooling strategies. We evaluate mean and last-token pooling after fine-tuning with Mistral-7B-Instruct-v0.1.
Categories $\longrightarrow$ Clas. Clus. P. Cls. Rera. Retr. STS Su. Average
# of datasets 12 11 3 4 15 10 1 56
Echo + last token pooling 77.43 46.32 87.34 58.14 55.52 82.56 30.73 64.68
Echo + mean pooling 77.00 44.94 87.73 58.30 55.11 82.52 29.46 64.22
Classical + last token pooling 76.57 45.78 86.37 56.71 54.87 82.03 31.02 63.98
Classical + mean pooling 76.26 42.68 86.31 57.58 53.75 81.53 30.19 62.96
Additional ablations.
We plot additional ablations, including ablating the role of instructions during training and evaluation, as well as providing an evaluation at step 280 (out of 720 total steps), which is approximately $40\%$ of the duration of training (Table 17). We note that echo embeddings still outperform classical embeddings in this setting.
Comparisons to the MTEB leaderboard.
We include some recent models on the MTEB leaderboard in Table 13.
Performance over training time.
All results.
NLI Retrieve semantically similar text
Squad Retrieve Wikipedia passages that answer the question
TriviaQA Retrieve Wikipedia passages that answer the question
Table 15: Instructions for finetuning datasets.
Training objective.
@@ Appendix F Additional Finetuning Results
Table 17: Additional ablations for finetuning. All ablations were performed using Mistral-7B as a backbone. We include ablations to compare the effect of using instructions, the effect of using the mean or last token, and the effect of using the last token at a different step.
Categories $\longrightarrow$ Clas. Clus. P. Cls. Rera. Retr. STS Su. Avg
# of datasets 12 11 3 4 15 10 1 56
Echo (w/ instruct., mean) 77.00 44.94 87.73 58.30 55.11 82.52 29.46 64.22
Echo (w/ instruct., last) 77.43 46.32 87.34 58.14 55.52 82.56 30.73 64.68
Classical (w/ instruct., mean) 76.26 42.68 86.31 57.58 53.75 81.53 30.19 62.96
Classical (w/ instruct., last) 76.57 45.78 86.37 56.71 54.87 82.03 31.02 63.98
Echo (w/out instruct., mean) 75.26 42.93 86.95 57.05 55.65 81.40 30.62 63.28
Echo (w/out instruct., last) 75.30 42.94 86.31 57.31 54.18 80.92 31.00 62.80
Classical (w/out instruct., mean) 75.23 41.79 85.24 56.31 53.24 80.97 30.64 62.19
Classical (w/out instruct., last) 75.01 42.70 85.69 56.64 53.29 80.92 30.91 62.37
Echo (w/ instruct., mean, step 280) 76.84 45.76 87.72 59.33 53.55 82.64 30.33 64.04
Echo (w/ instruct., last, step 280) 76.41 46.70 87.17 59.10 54.84 82.98 31.09 64.50
Classical (w/ instruct., mean, step 280) 76.18 42.99 85.44 57.63 53.96 82.53 29.94 63.19
Classical (w/ instruct., last, step 280) 76.54 46.22 86.70 57.79 53.73 82.22 30.13 63.87
Dataset Echo (last) Echo (mean) Classical (last) Classical (mean) Bidirectional (last)